OpenStudy (anonymous):

8 is 2/3 of what number

OpenStudy (unklerhaukus):

\[8=\frac 23\times n\] \[8\times\frac32=n\]

OpenStudy (ganpat):

x * 2/3 = 8 so x ??

OpenStudy (anonymous):

so what number is it

OpenStudy (anonymous):

x is not a number nor is n

OpenStudy (anonymous):

do the multiplication 8 x 3/2

OpenStudy (anonymous):

n= number

OpenStudy (unklerhaukus):

yeah n or x is the number you are looking for

OpenStudy (anonymous):

We have let the unknown number as n or x.. It depends upon you only that what number you want to let.. You can let it as b c d or anything you want..